OpenStudy (anonymous):

Find the future value of $12,430 borrowed at 6 5/8% simple interest for 810 days.

OpenStudy (tkhunny):

I = Prt -- Fix this in your brain. Future Value = P + I = P + Prt = P(1 + rt)